8. A particle executing a simple harmonic motion has the following an amplitude of 0.8 m and an frequency of 0.4 rads. (a) At what time, t, (after the particle started its motion) will the particle be at m? a) (b) What is the speed of the particle at this time?
Answer #1

A=0.8m

X=acoswt=0.8coswt=0.4

Coswt=0.5

Wt=60=pi/3

t=1/20s=0.05s

Velocity at t=1/20s

V=w(A2-x2)^1/2=14.5m/s substiute A=0.8 m and x=0.4m